# Scalable Multi-Agent Path Finding using Collision-Aware Dynamic Alert Mask and a Hybrid Execution Strategy

### v1

@@ Conclusion
A.2 Literature review
@@
S1 – Agent Planning: Distributed: In both PRIMAL and PRIMAL2, each agent independently plans its path using policies learned through a combination of RL and IL. These policies are trained to enable agents to navigate towards their goals based on local observations without centralized coordination (Sartoretti et al. 2019; Damani et al. 2021b).
@@
Table 2: Related Literature Categorization into – S1: Agent Planning; S2: Collision Detection; S3: Collision avoidance policy; S4: Agent Replanning (SB: Search Based; LB: Learning Based)
@@
Our* Decentralized Centralized Centralized Decentralized
@@ A.3 Training Methods
ResNetDQN is a residual‐network architecture for approximating the action‐value function $Q^{\pi}(s,a)$ in grid‐based environments trained using the DDQN algorithm (Van Hasselt, Guez, and Silver 2016). It combines a deep convolutional stem with residual blocks (He et al. 2016), early fusion of low‐dimensional features, hierarchical downsampling, and a late‐fusion MLP to output one Q‐value per action.
@@
Model Comparison
Table 3: Performance comparison of reinforcement learning methods (BFS-DQN, BFS-PPO, A*-DQN, A*-PPO) across different grid environments and agent counts. Results are averaged over 10 problem instances per configuration. Evaluated metrics include Success Rate (SR), Makespan (MS), Collisions (CO), and Time (T), with the Time metric averaged over all trials. Shows that the model trained using DDQN, selected for our main paper, yields better results over PPO model.
(a) 11x11 maze grid, $\sim$45% static obstacles (5–20 agents). Time in seconds.
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (s) ($\downarrow$)
Methods 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20
BFS-DDQN 100.00% 100.00% 98.00% 17.8 23.6 27.4 9 53 191 1.1 4.2 11.8
BFS-PPO 100.00% 94.00% 74.00% 22.1 28.6 28.6 9 48 125 4.9 16.5 31.8
A*-DDQN 100.00% 96.00% 82.00% 16.1 20.7 24.0 10 59 173 3.3 12.0 25.8
A*-PPO 100.00% 96.00% 74.00% 16.6 21.4 25.2 10 60 171 3.9 14.4 30.4
(b) 21x21 maze grid, $\sim$35% static obstacles (32, 64, and 96 agents). Time in minutes.
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
Methods 32 64 96 32 64 96 32 64 96 32 64 96
BFS-DDQN 100.00% 90.00% 0.00% 86.6 131.4 - 528 18909 - 2.7 23.9 33.5
BFS-PPO 100.00% 0.00% 0.00% 87.6 - - 463 - - 11.2 30.2 31.1
A*-DDQN 100.00% 80.00% 0.00% 41.4 72.0 - 457 11379 - 3.3 23.5 33.4
A*-PPO 100.00% 0.00% 0.00% 39.6 78.3 - 407 10761 - 6.2 27.0 30.2
(c) 25x25 warehouse grid, $\sim$24% static obstacles (32, 64, and 96 agents). Time in minutes.
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
Methods 32 64 96 32 64 96 32 64 96 32 64 96
BFS-DDQN 100.00% 100.00% 10.00% 94.3 125.9 96.0 237 6379 26553 5.2 28.0 59.5
BFS-PPO 100.00% 40.00% 0.00% 93.6 130.0 - 231 3740 - 15.2 57.3 60.2
A*-DDQN 100.00% 100.00% 60.00% 40.5 43.7 50.8 211 3129 16277 4.7 19.2 51.8
A*-PPO 100.00% 100.00% 40.00% 41.6 42.4 51.5 179 3117 17382 5.7 30.0 55.4
A.4 Additional Experiments
We present a model comparison trained using DDQN and PPO reinforcement learning algorithms, Table 3, in Section A.3. This section details further empirical evaluations and an ablation study of our framework.
An ablation study of different agent selection policies is detailed in Table 4. We have considered the information sharing setting/replanning strategy as Static & Dynamic Obstacle Replanning.
Furthermore, a comparison of replanning settings for static and dynamic obstacles is provided in Table 5, which reflects different information sharing settings as detailed in our main paper. Here, we considered the agent selection policy as Fewest Future Collisions (FFC).
We also present extended analyses on MovingAI lab’s MAPF benchmark dataset (Stern et al. 2019b) – random map $32\times 32-10$ environment with various agent counts [8, 16, 32, 64, 96], results presented in Section A.4.
Agent Selection Policy Comparison
Table 4: Performance comparison of Agent Selection Policies (Random, Farthest, Fewest Future Collisions - FFC) for Alert-BFS and Alert-A* methods across different grid environments and agent counts. Metrics include Success Rate (SR), Makespan (MS), Collisions (CO), and Time (T). Results are averaged over 10 problem instances per configuration. Best performance across policies for each method and agent group is bolded. Using Static & Dynamic obstacle replanning. Shows that, among the choices [random, farthest & fewest future collisions (FFC)], the FFC policy, selected for our main paper, yields best results across the different map configurations.
(a) 11x11 maze, $\sim$45% static obstacles (5–20 agents). Time in seconds (s).
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (s) ($\downarrow$)
Methods 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20
Policy: Random Agent Selection
Alert-BFS 98.33% 98.00% 82.00% 20.5 24.6 32.7 20 108 517 2.0 5.8 22.6
Alert-A* 98.33% 96.00% 50.00% 17.8 23.6 24.3 16 122 259 4.3 15.7 35.6
Policy: Farthest Agent Selection
Alert-BFS 100.00% 94.00% 90.00% 19.8 24.6 30.3 14 91 387 1.2 6.9 17.8
Alert-A* 98.33% 84.00% 62.00% 18.9 23.3 26.5 14 96 224 4.6 18.3 35.1
Policy: Fewest Future Collisions (FFC)
Alert-BFS 100.00% 100.00% 98.00% 17.8 23.6 27.4 9 53 191 1.1 4.2 11.8
Alert-A* 100.00% 96.00% 82.00% 16.1 20.7 24.0 10 59 173 3.3 12.0 25.8
(b) 21x21 maze grid, $\sim$35% static obstacles (32, 64, and 96 agents). Time in minutes (min).
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
Methods 32 64 96 32 64 96 32 64 96 32 64 96
Policy: Random Agent Selection
Alert-BFS 100.00% 0.00% 0.00% 99.2 - - 1670 - - 4.1 30.1 30.4
Alert-A* 100.00% 0.00% 0.00% 48.8 - - 931 - - 6.1 30.2 30.2
Policy: Farthest Agent Selection
Alert-BFS 100.00% 0.00% 0.00% 93.5 - - 1216 - - 3.1 30.2 30.2
Alert-A* 100.00% 10.00% 0.00% 47.5 64.0 - 697 14654 - 5.3 29.8 30.3
Policy: Fewest Future Collisions (FFC)
Alert-BFS 100.00% 90.00% 0.00% 86.6 131.4 - 528 18909 - 2.7 23.9 30.5
Alert-A* 100.00% 80.00% 0.00% 41.4 72.0 - 457 11379 - 3.3 23.5 30.4
(c) 25x25 warehouse grid, $\sim$24% static obstacles (32, 64, and 96 agents). Time in minutes (min).
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
Methods 32 64 96 32 64 96 32 64 96 32 64 96
Policy: Random Agent Selection
Alert-BFS 100.00% 20.00% 0.00% 119.0 139.0 - 722 23853 - 12.4 57.9 60.1
Alert-A* 100.00% 100.00% 0.00% 42.4 54.6 - 302 5788 - 6.0 37.9 60.1
Policy: Farthest Agent Selection
Alert-BFS 100.00% 60.00% 0.00% 83.4 126.0 - 455 9836 - 8.5 47.6 60.8
Alert-A* 100.00% 100.00% 0.00% 45.0 59.4 - 268 4644 - 6.4 34.9 60.1
Policy: Fewest Future Collisions (FFC)
Alert-BFS 100.00% 100.00% 10.00% 94.3 125.9 96.0 237 6379 26553 5.2 28.0 59.5
Alert-A* 100.00% 100.00% 60.00% 40.5 43.7 50.8 211 3129 16277 4.7 19.2 51.8
Inter-agent Information Sharing Comparison
Table 5: Comparison of replanning settings (Only Static, Only Dynamic, Static & Dynamic Obstacles) for Alert-BFS and Alert-A* methods across different grid environments and agent counts. Metrics include Success Rate (SR), Makespan (MS), Collisions (CO), and Time (T). Results are averaged over 10 problem instances per configuration. Best performance across settings for each method and agent group is bolded. Using FFC agent selection policy. Shows that, among the replanning strategies, Static + Dynamic obstacle replanning strategy (selected for our main paper), yields best results across the different map configurations.
(a) 11x11 maze grid, $\sim$45% static obstacles (5–20 agents). Time in seconds (s).
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (s) ($\downarrow$)
Methods 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20
Setting 1: Only Static Obstacle Replanning (No inter-agent information)
Alert-BFS 95.00% 92.00% 84.00% 17.5 23.0 25.4 8 51 205 1.4 4.6 14.2
Alert-A* 95.00% 90.00% 78.00% 15.6 20.3 23.7 8 55 177 3.2 11.6 26.0
Setting 2: Only Dynamic Obstacle Replanning
Alert-BFS 63.33% 24.00% 8.00% 19.3 21.0 26.8 16 89 171 4.9 17.9 34.7
Alert-A* 50.00% 12.00% 2.00% 14.0 13.8 15.0 6 25 49 22.9 41.4 52.1
Setting 3: Static & Dynamic Obstacle Replanning
Alert-BFS 100.00% 100.00% 98.00% 17.8 23.6 27.4 9 53 191 1.1 4.2 11.8
Alert-A* 100.00% 96.00% 82.00% 16.1 20.7 24.0 10 59 173 3.3 12.0 25.8
(b) 21x21 maze grid, $\sim$35% static obstacles (32, 64, and 96 agents). Time in minutes (min).
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
Methods 32 64 96 32 64 96 32 64 96 32 64 96
Setting 1: Only Static Obstacle Replanning (No inter-agent information)
Alert-BFS 100.00% 10.00% 0.00% 86.6 140.0 - 528 12476 - 5.8 30.0 30.0
Alert-A* 100.00% 20.00% 0.00% 41.4 59.0 - 457 9310 - 6.7 30.0 30.0
Setting 2: Only Dynamic Obstacle Replanning
Alert-BFS 0.00% 0.00% 0.00% - - - - - - 30.0 30.0 30.0
Alert-A* 0.00% 0.00% 0.00% - - - - - - 30.0 30.0 30.0
Setting 3: Static & Dynamic Obstacle Replanning
Alert-BFS 100.00% 90.00% 0.00% 86.6 131.4 - 528 18909 - 2.7 23.9 30.0
Alert-A* 100.00% 80.00% 0.00% 41.4 72.0 - 457 11379 - 3.3 23.5 30.0
(c) 25x25 warehouse grid, $\sim$24% static obstacles (32, 64, and 96 agents). Time in minutes (min).
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
Methods 32 64 96 32 64 96 32 64 96 32 64 96
Setting 1: Only Static Obstacle Replanning (No inter-agent information)
Alert-BFS 100.00% 100.00% 0.00% 95.8 117.4 - 230 6097 - 10.0 40.0 60.2
Alert-A* 100.00% 100.00% 60.00% 42.2 41.6 56.0 204 3258 16694 8.9 32.3 53.3
Setting 2: Only Dynamic Obstacle Replanning
Alert-BFS 0.00% 0.00% 0.00% - - - - - - 61.7 62.4 60.6
Alert-A* 0.00% 0.00% 0.00% - - - - - - 58.6 61.6 61.8
Setting 3: Static & Dynamic Obstacle Replanning
Alert-BFS 100.00% 100.00% 10.00% 94.3 125.9 96.0 237 6379 26553 5.2 28.0 59.5
Alert-A* 100.00% 100.00% 60.00% 40.5 43.7 50.8 211 3129 16277 4.7 19.2 51.8
Benchmark Dataset Evaluation
Map Configuration: $32\times 32-10$ Random Map
The random map $32\times 32-10$ is obtained from the Moving AI Lab’s MAPF benchmark dataset (Stern et al. 2019b). Agent configurations tested include scenarios with 8, 16, 32, 64, and 96 agents. Each configuration was tested across 10 different problem instances (unique start and goal locations for all agents). The detailed performance metrics, including success rates, makespan, collisions, and computation time, are presented in Table 6.
Table 6: Comparison of performance metrics for different methods on a $32\times 32$ random map. Results are averaged over 10 problem instances per configuration and presented in two subtables based on agent counts. Key metrics include Success Rate (SR), Makespan (MS), Collisions (CO), and Time (T(min)), with Time averaged over all trials. For learning-based methods like PRIMAL and SCRIMP, the average steps taken (episode length) are presented in brackets along with average wall-clock time taken. For learning based methods, the environment step limit is set to 512. Best performing values for each metric per agent count are highlighted in bold.
(a) 32x32 random map (8, 16, and 32 agents). The time limit per problem instance is 1 hour. For learning based methods, the environment step limit is set to 512.
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
Methods 8 16 32 8 16 32 8 16 32 8 16 32
Alert-BFS 100.00% 100.00% 100.00% 62.5 76.2 104.9 2 9 64 0.2 0.6 2.4
Alert-A* 100.00% 100.00% 100.00% 42.7 45.5 48.1 2 8 68 0.2 0.9 2.5
CBS 100.00% 100.00% 90.00% 40.2 43.6 47.0 1 6 35079 0.01 0.01 13.0
ICBS 100.00% 100.00% 90.00% 40.2 43.6 47.0 1 6 15438 0.01 0.01 10.7
PRIMAL 0.00% 0.00% 0.00% - - - - - - 0.1 (512) 0.3 (512) 1.0 (512)
SCRIMP 100.00% 100.00% 100.00% 39.4 43.3 47.7 0 0 0 0.01 (39) 0.01 (43) 0.03 (48)
(b) 32x32 random map (64 and 96 agents). The time limit per problem instance is 1 hour. For learning based methods, the environment step limit is set to 512.
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
Methods 64 96 64 96 64 96 64 96
Alert-BFS 100.00% 100.00% 136.5 154.4 1194 7079 12.9 33.8
Alert-A* 100.00% 100.00% 53.2 52.0 1158 5912 13.2 30.7
CBS 0.00% 0.00% - - - - 60.1 60.1
ICBS 0.00% 0.00% - - - - 60.3 60.2
PRIMAL 0.00% 0.00% - - - - 3.6 (512) 8.4 (512)
SCRIMP 100.00% 90.00% 52.1 55.8 2 3 0.2 (52) 0.5 (101)

### v2

@@ Conclusion
A.2 Properties of Tiered Replanning Strategies
In this section, we summarize the soundness and (local) completeness properties of the tiered replanning operators introduced in S4 (Replanning). We assume a standard discrete-time MAPF model on a finite grid graph $G=(V,E)$ with a finite set of agents $\{a_{1},\dots,a_{N}\}$. Time is discrete, $t=0,1,\dots,T_{\max}$; each agent can either move along an edge $(u,v)\in E$ or wait at its current vertex. A plan for agent $a_{i}$ is a finite sequence of vertices
$\pi_{i}=(v_{i}(0),v_{i}(1),\dots,v_{i}(T_{i})),$
where $(v_{i}(t),v_{i}(t+1))$ is either an edge in $E$ or a wait, with $v_{i}(0)$ and $v_{i}(T_{i})$ denoting its fixed start and goal positions, respectively. A joint plan is the collection $\Pi=\{\pi_{1},\dots,\pi_{N}\}$.
We adopt the standard MAPF collision model: a joint plan $\Pi$ is collision-free if it contains no vertex collision
$\exists i\neq j,\ \exists t:\ v_{i}(t)=v_{j}(t),$
and no edge collision
$\exists i\neq j,\ \exists t:\ v_{i}(t)=v_{j}(t+1)\ \wedge\ v_{i}(t+1)=v_{j}(t).$
Throughout this discussion, we assume that each tier uses an underlying search routine (single-agent planner or joint planner) that is complete on its induced finite state-time graph (e.g., BFS or A* with an admissible, consistent heuristic), and that any candidate repair is re-simulated against the other agents before being committed.
Soundness.
We first formalize soundness of a replanning operator.
Definition 1 (Soundness)
A replanning operator $\mathcal{R}$ is sound if, whenever it returns a modified joint plan $\Pi^{\prime}$, the resulting joint plan is collision-free under the given collision model.
In our framework, S4.1–S4.4 operate under constraints provided by the control module (S3) and are only accepted if a global collision check succeeds.
Lemma 1 (Soundness of S4.1–S4.4)
Assume that each replanning operator (yield-based, bounded static, bounded dynamic, and local joint planning) only considers legal moves (graph edges or waits), obeys all constraints imposed by S3 (forbidden cells, reservations, and joint consistency), and that any candidate repaired plan is re-simulated against all other agents and discarded if any collision is detected. Then every committed repair produced by S4.1–S4.4 is collision-free; in particular, each tier is sound.
Proof 1 (Proof sketch)
Each operator constructs candidate paths by exploring a constrained state-time graph in which illegal moves (violating static constraints, dynamic reservations, or internal joint consistency) are excluded. By construction, no candidate produced by the underlying planner violates these constraints. Before committing any repair, the resulting joint plan $\Pi^{\prime}$ is simulated and checked for vertex and edge collisions against all other agents. If a collision is found, the candidate is rejected. Hence any repair that is actually committed has passed both constraint enforcement and a global collision check, and is therefore collision-free.
Yield-based local coordination (S4.1).
The yield operator chooses a single agent (possibly anchored at its start or goal) and constructs a short detour to a nearby “parking” cell, waits there while other agents traverse the conflicted region, and then rejoins the original plan. The search for a parking cell is restricted to a bounded local neighborhood and does not modify other agents’ plans.
Lemma 2 (Soundness and incompleteness of S4.1)
The yield-based operator S4.1 is sound but not complete even for resolving a single conflict: there exist MAPF instances in which a collision can be resolved by modifying agents’ plans, but no admissible yield maneuver exists within the bounded neighborhood and single-agent restriction imposed by S4.1.
Proof 2 (Proof sketch)
Soundness follows directly from Lemma 1: any candidate yield path is simulated and only committed if no collisions remain.
To see incompleteness, consider a long, narrow corridor of width one, and two agents at opposite ends in the middle of the corridor, wishing to swap positions. A known solution exists (e.g., coordinating their timing so one waits at the opening while the other passes), but under a strict yield model that only searches for a parking cell in a small neighborhood around the collision, it is impossible to find a nearby parking position that both satisfies the local constraints and removes the conflict. More generally, whenever resolving a conflict requires modifying the timing or paths of multiple agents together, or using parking locations outside the bounded neighborhood, S4.1 may fail to find a repair even though a global solution exists. Thus, S4.1 is sound but incomplete.
Bounded static replanning (S4.2).
In bounded static replanning, the control module S3 selects a single agent, fixes all other agents’ plans, and designates a time window $\{t_{\mathrm{start}},t_{\mathrm{end}}\}$ along the agent’s current plan. A static constraint set $\Delta_{c}\subseteq V$ encodes cells that must be avoided (e.g., conflict cells), and the planner recomputes only the affected suffix of the agent’s trajectory while treating $\Delta_{c}$ as forbidden and all other agents’ plans as fixed reservations.
This induced subproblem can be modeled as single-agent planning on a finite state-time graph whose nodes are $(v,t)$ and whose edges correspond to legal moves that respect static constraints and avoid collisions with fixed agents.
Lemma 3 (Local completeness of S4.2)
Consider the state-time graph induced for S4.2 over a finite horizon, with all vertices $(v,t)$ such that $v\in\Delta_{c}$ in the relevant window removed, and all nodes and edges that would collide with fixed agents pruned. Assume the single-agent planner used in S4.2 is complete on this graph (e.g., BFS or A* with an admissible, consistent heuristic). If there exists a collision-free path for the selected agent from its state at $t_{\mathrm{start}}$ to its target (goal or intermediate waypoint) within this graph, then S4.2 will find such a path and return a valid repair.
Proof 3 (Proof sketch)
The state-time graph is finite, as both the underlying grid and the time horizon are finite. Every collision-free path that respects static constraints and avoids fixed agents corresponds to a path in this pruned graph. A complete search algorithm on a finite graph is guaranteed to find a path to the target node if one exists. Therefore, under the assumed planner, S4.2 is complete for its induced subproblem.
Bounded dynamic replanning (S4.3).
Bounded dynamic replanning extends S4.2 by treating portions of other agents’ plans as time-varying obstacles (reservations) over a finite horizon around the conflict. In the corresponding state-time graph, nodes and edges that would cause vertex or edge collisions with these reservations are removed.
Lemma 4 (Local completeness of S4.3)
Under the same assumptions as Lemma 3, but with additional time-indexed reservations derived from other agents’ plans, bounded dynamic replanning (S4.3) is complete for its induced single-agent subproblem: if there exists a collision-free path that respects both static constraints and dynamic reservations in the finite horizon, the underlying complete planner will find it.
Proof 4 (Proof sketch)
The state-time graph with dynamic reservations remains finite; reservations simply prune additional nodes and edges corresponding to disallowed states and transitions. Any path that respects all reservations corresponds to a path in this pruned graph. A complete search will discover such a path if it exists. Hence S4.3 is locally complete for its induced single-agent subproblem.
Local joint planning (S4.4).
When single-agent repairs are insufficient, S4.4 jointly replans a small subset of agents $S\subseteq\{1,\dots,N\}$ over a bounded time window around the conflict, treating all other agents as fixed reservations. The induced local problem can be modeled as a joint state-time graph whose nodes are $(\mathbf{v}(t),t)$, where $\mathbf{v}(t)=(v_{k}(t))_{k\in S}$ is a joint configuration with no internal vertex collisions and no collisions with reserved agents.
Lemma 5 (Local completeness of S4.4)
Consider the joint state-time graph for a fixed subset of agents $S$ and a finite horizon $\{t_{\mathrm{start}},t_{\mathrm{end}}\}$, where nodes correspond to collision-free joint configurations of $S$ and edges correspond to legal joint moves (Cartesian products of individual moves) that do not introduce collisions among $S$ or with reserved agents. Assume that S4.4 runs a complete joint planner (e.g., joint A* with an admissible heuristic) on this graph. If there exists a collision-free joint path from the initial configuration of $S$ at $t_{\mathrm{start}}$ to the designated joint target region by $t_{\mathrm{end}}$, then S4.4 will find such a joint path.
Proof 5 (Proof sketch)
The joint state-time graph is finite: the number of joint configurations is bounded by $|V|^{|S|}$ per time step, and the time horizon is finite. Edges are finite and defined by admissible joint actions. A complete search algorithm on this finite graph is guaranteed to find a path to a reachable goal region. Any valid joint plan corresponds to such a path. Therefore S4.4 is complete for its induced joint subproblem.
Discussion and overall behavior.
The above lemmas show that each tier is sound and that the static, dynamic, and joint operators (S4.2–S4.4) are complete for their respective induced subproblems under the assumption of complete search over the chosen finite horizon. Yield-based coordination (S4.1) is sound but intentionally incomplete. In the full algorithm, these tiers are combined with a deferral mechanism: if all tiers fail for an agent within the current attempt and bounded parameters, the agent is temporarily deferred (parked and excluded from further replanning), the remaining agents are resolved, and the deferred agents are reintroduced in a second phase with relaxed limits.
With fixed bounds on time windows, joint subset sizes, and search effort, the overall tiered strategy is best viewed as a sound, information-efficient repair framework with local completeness guarantees, rather than a globally complete MAPF algorithm. In principle, completeness for the full MAPF instance could be recovered by allowing the temporal windows and sizes of jointly replanned subsets to grow as needed and by repeatedly applying the tiers (including deferral) until all conflicts are resolved; however, this configuration would come at a substantially higher computational cost and is not pursued in our current implementation.
Figure 4: Five map configuration, initial state and goal state, used for testing our IC–MAPF algorithm on TurtleBot4s’
Figure 5: Indoor 6x6 grid setup with TurtleBot4’
A.3 Hardware Results: TurtleBot4
We evaluated IC–MAPF on five TurtleBot4 robots deployed in a $6\times 6$ indoor grid. The centralized coordinator module (S3) ran on a central host PC, while each robot also executed its own local planning (S1) and replanning procedures (S4) under isolated ROS 2 namespaces. Robots communicated only through a shared synchronization topic (/sync_barrier) and operated under a ROS 2 Discovery Server setup, ensuring that no robot was configured to discover or access any other robot’s topics or services; the only cross-namespace visibility permitted was the synchronization channel. No onboard perception (LiDAR, RGB-D, or OAK-D) was used; instead, robots followed their assigned plans using open-loop odometry control. All remaining ROS interfaces, including odometry and velocity commands, remained fully namespace-isolated, ensuring minimal inter-agent information exchange.
At the start of each run, robots reset their poses via the /reset_pose service and waited for odometry stabilization. Execution proceeded in strict lockstep: at every step, a robot (i) rotated to the required heading using closed-loop angular control, and (ii) moved forward by $0.45$ m using odometry-based distance tracking. After completing its action, each robot published a synchronization message and waited until all robots reached the same step, ensuring coordinated, collision-free execution across the team. We evaluated five MAPF problems, each repeated five times, and logged communication statistics, execution times, and stepwise deviations from planned poses. Communication and execution metrics are reported in Table 4, while execution-quality statistics are summarized in Table 5.
These experiments demonstrate that IC–MAPF can be reliably deployed on real robots with extremely lightweight sensing and communication requirements. In contrast, learning-based MAPF methods that depend on field-of-view observations would require additional sensors, onboard neural inference, higher energy consumption, and greater bandwidth usage. IC–MAPF’s ability to operate using only minimal synchronization messages highlights its practicality for resource-constrained, scalable, and privacy-preserving multi-robot systems.
Map Plan. Msgs IU Exec. Msgs Exec. Time (s)
Map 1 6 48 55 110.3
Map 2 16 15 75 162.6
Map 3 8 21 65 129.5
Map 4 8 14 65 145.7
Map 5 10 54 95 260.5
Table 4: Communication and execution-related statistics for five TurtleBot4 trials.
Map MS Pos. Err. (cm) Head. Err. ()
Map 1 11 2.08 1.97
Map 2 15 2.63 2.42
Map 3 13 2.14 2.09
Map 4 13 2.12 2.01
Map 5 19 2.58 2.54
Table 5: Execution quality measures: makespan, average positional deviation, and heading deviation for five TurtleBot4 runs. Harder problems (higher MS) show slightly larger systematic drift.
A.4 Literature review
@@
S1 – Agent Planning: Distributed: In both PRIMAL and PRIMAL2, each agent independently plans its path using policies learned through a combination of RL and IL. These policies are trained to enable agents to navigate towards their goals based on local observations without centralized coordination (Sartoretti and others 2019; Damani et al. 2021).
@@
Table 6: Related Literature Categorization into – S1: Agent Planning; S2: Collision Detection; S3: Collision avoidance policy; S4: Agent Replanning (SB: Search Based; LB: Learning Based). Our approach is IC–MAPF with S4 variants 4.0-4.3.
@@
Our: IC–MAPF Decentralized Centralized Centralized Distributed (4.0 - 4.3)
@@ A.5 Training Methods
ResNetDQN is a residual‐network architecture for approximating the action‐value function $Q^{\pi}(s,a)$ in grid‐based environments trained using the DDQN algorithm (Van Hasselt et al. 2016). It combines a deep convolutional stem with residual blocks (He et al. 2016), early fusion of low‐dimensional features, hierarchical downsampling, and a late‐fusion MLP to output one Q‐value per action.